The crankpin of an engine sustains a maximum load of 35 kN due to steam pressure. If the allowable

bearing pressure is 7 N/mm2


, find the dimensions of the pin. Assume the length of the pin equal to 1.2

times the diameter of the pin.


1
Expert's answer
2021-10-29T02:33:17-0400

diameter of the pin. d

"\sigma = \frac{W}{A} = \frac{W}{\frac{\pi}{4}d^2}"


"d^2 = \frac{35000}{\frac{\pi}{4}\times 7 }"

"d = 79.78 mm \iff80 mm"

diameter of the pin. d = 80 mm

Length of Pin, "L=1.2d =1.2\times 80 =96mm"